# CacheFlow: Fast Human Motion Prediction by Cached Normalizing Flow

### v1

@@ 4.3 Conditional Inference by CacheFlow
In previous methods, conditional human motion prediction typically requires a conditional generative model. For instance, it is a conditional flow-based or diffusion model. These models usually have poor time efficiency due to delicate but heavy architecture. Instead, inspired by Equation 9, we can reuse the results of unconditional inverse transformation as triplets $t_{k}=\{\bm{z}_{k},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z}_{k})|^{-1},\bm{x}_{k}\}$. Thus, we can perform conditional inference by only evaluating a conditional base distribution $q(\bm{z}|\bm{c})$. We model this conditional base distribution by a learnable model, thus we denote it as $q_{\phi}(\bm{z}|\bm{c})$. This model can be very lightweight since the unconditional transformation $f_{\theta}$ gives enough expressivity. $q_{\phi}(\bm{z}|\bm{c})$ runs much faster than a typical conditional generative model for human motion prediction. This part is illustrated in Figure 2(b).
In our implementation, $q_{\phi}(\bm{z}|\bm{c})$ is constructed as a parametrized Gaussian mixture $\{\mathcal{N}(\mu_{m}(\bm{c}),\sigma^{2}_{m}(\bm{c}))\}$, with $M$ mixture weights $w_{m}(\bm{c})$, such that $\sum_{m=1}^{M}w_{m}=1$. Each $\mu_{m}$ and $\sigma_{m}$ are regressed based on the feature of past motion $\bm{c}$. We use a lightweight single-layer RNN for regression to determine the GMM composition. Although the unconditional flow-based model $f_{\theta}$ and the conditional base density $q_{\phi}$ can be trained separately, we found that jointly training $f_{\theta}$ and $q_{\phi}$ improves model performance. We train the joint model by summation of log-likelihood for $q_{\phi}$ and flow matching for $f_{\theta}$ as explained in Equation 6 as follows:
@@
$\begin{split}k^{*}&={\rm argmin}_{k}||\bm{z}_{k}-\bm{z}||,\\ {\rm s.t.}\quad\{t_{k}\ &=\{\bm{z}_{k},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z}_{k})|^{-1},\bm{x}_{k}\},\quad\bm{z}\sim q_{\phi}(\bm{z}|\bm{c})\},\end{split}$ (13)
@@ 5 Experimental Evaluation
Datasets. We evaluate our CacheFlow on Human3.6M [25] and AMASS [42]. Human3.6M contains 3.6 million frames of human motion sequences. Human motions of 11 subjects performing 15 actions are recorded at 50 Hz. We follow the setting including the dataset split, the 16-joints pose skeleton definition, and lengths of past and future motions proposed by previous works [47, 38, 71, 54]. The training and test sets of Human3.6M are subjects [S1,S5,S6,S7,S8] and [S9,S11], respectively. The past motion and future motions contain 25 frames (0.5 sec) and 100 frames (2.0 sec). AMASS unifies 24 different human motion datasets including HumanEva-I [59] with the SMPL [41] pose representation. AMASS contains 9M frames at 60 Hz in total. As a multi-dataset collection of AMASS, one can perform a cross-dataset evaluation. We follow the evaluation protocol proposed by BeLFusion [4] for fair comparison, as predicting future 120 frames (2.0 sec) with 30 frames observation (0.5 sec) with downsampling to 60 Hz.
Metrics. We use the evaluation metrics to measure diversity and accuracy. 50 sampled predictions are evaluated with the following metrics: Average Pairwise Distance (APD) [3] evaluates sample diversity. It calculates the mean $l_{2}$ distance between all predicted motions. Average and Final Displacement Error (ADE, FDE) [2, 32, 22] evaluate accuracy. They calculate the average and final-frame $l_{2}$ distances between the ground truth motion and closest prediction in the 50 set. Multimodal ADE and FDE (MMADE, MMFDE) [72] also evaluate accuracy in a similar way to ADE and FDE. However, they are calculated over multimodal ground truths selected by grouping similar motions.
@@
Implementation Details. Our method is based on a latent flow-based model. We utilize a Variational Autoencoder (VAE) to obtain a latent representation. Specifically, we employ the Behavioral Latent Space (BLS) [4] as a VAE to achieve a compact latent representation. BLS ensures smoothness of predicted motions and consistency between the end of the past motion and the start of the predicted motion. Additionally, we compress this representation using linear factorization [68]. The dimensionality of the VAE latent space is 128, which we further reduce to 8 dimensions through linear factorization. We trained the unconditional flow-based model on this 8-dimensional space. The unconditional flow-based model $f_{\theta}$ is a continuous normalizing flow (CNF) model, with its vector field regressed by a U-Net architecture. The conditional base density $q_{\phi}$, as well as the VAE encoder and decoder, are implemented as one-layer Recurrent Neural Networks (RNNs). We used a Gaussian mixture model with $M=50$ modes to model the conditional base density $q_{\phi}$. We precomputed and collected triplets $t_{k}=\{\bm{z}_{k},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z}_{k})|^{-1},\bm{x}_{k}\}$ using all training samples of each dataset. All experiments, including inference time measuring, were carried out using a single NVIDIA A100 GPU. We used a batch size 64 and the Adam optimizer with a learning rate of $5\times 10^{-4}$.
@@ 5.1 Quantitative Evaluation
CoMusion and SLD were successful in predicting motions that are closer to the ground truth than CacheFlow; however, their inference times of 167 and 375 milliseconds are too long for the intended 2000 ms prediction horizon. As a result, over 8% of the first prediction sequence is rendered useless once the prediction is finalized. Therefore, it is difficult to use these methods with slow inference in real-time applications. Although our primary goal is to estimate the density, CacheFlow achieves comparable performances with a 1.3 millisecond inference time. Our method achieves around 4$\times$ faster than the fastest VAE method, GSPS, and 30$\times$ faster than the fastest diffusion-based method, BeLFusion-D. The inference of our method is fast enough (1.3ms for future 2000ms) and applicable for real-time applications. This inference speed is because the inference of the unconditional flow-based model $f_{\theta}$ is precomputed. We only need to evaluate the lightweight conditional base density $q_{\phi}$ at inference. Although our conditional base density $q_{\phi}$ is just a Gaussian mixture with low expressive power, our method achieves high accuracy since the precomputed unconditional flow-based model $f_{\theta}$ gives $q_{\phi}$ much complexity with almost no overhead in inference.
@@
The density estimation accuracy of each method is compared between CacheFlow and the state-of-the-art methods. The three state-of-the-art methods BeLFusion [4], CoMusion [61], and SLD [68] are selected. CoMusion and SLD were selected since they outperform our method in benchmarks of stochastic human motion prediction. We also include BeLFusion to compare CacheFlow with the method with latent space. We applied KDE to these previous methods since they only sample a set of predictions and cannot estimate density. While we evaluated 50 and 1000 samples for KDE on BeLFusion and CoMusion, SLD only allows 50 samples due to the fixed number of anchors corresponding to predictions. We measured the inference time of each method to estimate the density of ten thousand future motions from the past motion input.
The quantitative comparisons over the multimodal ground truth log probability are shown in Table 2. All previous methods suffer from slow inference of their own and KDE on high-dimensional motion data. Their inference time exceeded the prediction horizon of 2000ms in the future. Therefore, they cannot estimate density in real-time. In contrast, our method achieves better estimation accuracy in less than one millisecond. This indicates that CacheFlow has strong discriminative ability to list up possible future motions required for safety assurance. Our method is even faster only on the density estimation (0.5ms) than the inference time reported in Table 1 (1.3ms). This is because we don’t need any extra sampling operation in the density estimation.
@@ 5.2 Qualitative Comparison of Predicted Motions
To visually evaluate CacheFlow, we conducted a qualitative comparison of methods on the AMASS dataset, as shown in Figure 3. We visualized the end poses of 10 samples from each method alongside the end poses of past motions and the ground truth future motions. The sitting or lying poses were translated to the ground plane, as the global translation is not modeled in human motion prediction. The 10 pose samples are arranged from the closest to the farthest from the ground truth pose based on joint rotations.
@@
Our observations indicate that CacheFlow predicts realistic poses. The closest poses to the ground truths also demonstrate that the accuracy of CacheFlow is comparable to CoMusion, as reflected in the ADE and FDE metrics listed in Table 1. Notably, our method is computationally efficient, operating 100 times faster than the fastest CoMusion. In summary, CacheFlow effectively delivers realistic and accurate predictions.
@@ 6 Concluding Remarks
Appendix A Implementation Details of Kernel Density Estimation
@@
Appendix B Ablation Study
[…]
We conducted an ablation study to investigate how each component affects the performance of our CacheFlow. We ablate five components: (1) dimensionality reduction via linear factorization on VAE, (2) the unconditional flow-based model $f_{\theta}$, (3) joint learning of the conditional base density $q_{\phi}$ and unconditional flow-based model $f_{\theta}$, (4) dataset for precomputation, (5) the sampling method for metrics over a fixed number of predictions. Ablation results on the Human3.6M dataset are summarized in Table 3.
The Unconditional Flow-based Model. We ablate this flow-based model $f_{\theta}$ to confirm it improves the conditional base density $q_{\phi}$ by adding complexity. As shown in Table 3, we observe a notable performance drop without the flow-based model. Therefore, our unconditional flow-based model $f_{\theta}$ complements conditional base density $q_{\phi}$ to estimate complex density distribution over human motions.
Dataset for Precomputation. We propose the precomputation over the training split. Specifically, we apply inverse transform $\bm{z}=f_{\theta}(\bm{x})$ to ground truth future motions in the training split. However, we may precompute infinite precomputation samples. For example, we can sample $\bm{z}\sim q_{\phi}(\bm{z}|\bm{c})$ and obtain $\bm{x}$ by forward transform $\bm{x}=f_{\theta}(\bm{z})$. As shown in the ablation, precomputation on the training split outperforms one on the base density since we can regularize the prediction to be legitimate human motions using the training split.
Sampling Method. We propose the nearest neighbor sampling from the precomputation set as described in Section 4.3. Lastly, we ablate this sampling to evaluate its performance gain. We experimented with two sampling method alternatives: random sampling and most likely sampling. Precomputed motion features $\bm{x}_{k^{*}}$ are uniformly selected as predictions with random sampling. Most likely sampling selects motion features $\bm{x}_{k^{*}}$ with the highest probabilities $k^{*}={\rm argmax}_{k}p(\bm{x}_{k}|\bm{c})$. We found that the large and little performance drops with random and most likely sampling respectively. This random sampling is worse due to the independence from the past motions $\bm{c}$. The most likely method underperforms due to less diverse samples. It cannot select a motion feature set with diversity because all selected features are often located in one peak of the estimated density. Since ADE and FDE are best-of-many metrics, this less diversity leads to worse performance. In contrast, our sampling method is superior to others. Our sampling incorporates past motions and achieves good diversity by simulating sampling from the estimated density $p(\bm{x}|\bm{c})$.

### v2

@@ 4.3 Conditional Inference by CacheFlow
In previous methods, conditional human motion prediction typically requires a conditional generative model. For instance, it is a conditional flow-based or diffusion model. These models usually have poor time efficiency due to delicate but heavy architecture. Instead, inspired by Eq.˜9, we can reuse the results of unconditional inverse transformation as triplets $t_{k}=\{\bm{z}_{k},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z}_{k})|^{-1},\bm{x}_{k}\}$. Thus, we can perform conditional inference by only evaluating a conditional base distribution $q(\bm{z}|\bm{c})$. We model this conditional base distribution by a learnable model, thus we denote it as $q_{\phi}(\bm{z}|\bm{c})$. This model can be very lightweight since the unconditional transformation $f_{\theta}$ gives enough expressivity. $q_{\phi}(\bm{z}|\bm{c})$ runs much faster than a typical conditional generative model for human motion prediction. This part is illustrated in Fig.˜2(b).
In our implementation, $q_{\phi}(\bm{z}|\bm{c})$ is constructed as a parametrized Gaussian mixture $\{\mathcal{N}(\mu_{m}(\bm{c}),\sigma^{2}_{m}(\bm{c}))\}$, with $M$ mixture weights $w_{m}(\bm{c})$, such that $\sum_{m=1}^{M}w_{m}=1$. Each $\mu_{m}$ and $\sigma_{m}$ are regressed based on the past motion $\bm{c}$. We use a lightweight single-layer RNN for regression to determine the GMM composition. This RNN takes the past motion directly as input, and regress the $w_{m},\mu_{m},\sigma^{2}_{m}$. Therefore, $\phi$ means the parameter of this RNN in our method. Although the unconditional flow-based model $f_{\theta}$ and the conditional base density $q_{\phi}$ can be trained separately, we found that jointly training $f_{\theta}$ and $q_{\phi}$ improves model performance. We train the joint model by summation of log-likelihood for $q_{\phi}$ and flow matching for $f_{\theta}$ as explained in Eq.˜6 as follows:
@@
[…]
@@ 5 Experimental Evaluation
Datasets. We evaluate our CacheFlow on Human3.6M (ionescu2013h36m) and AMASS (mahmood2019amass). Human3.6M contains 3.6 million frames of human motion sequences. Human motions of 11 subjects performing 15 actions are recorded at 50 Hz. We follow the setting including the dataset split, the 16-joints pose skeleton definition, and lengths of past and future motions proposed by previous works (martinez2017simple; luvizon20182d; yang20183d; pavllo20193d). The training and test sets of Human3.6M are subjects [S1,S5,S6,S7,S8] and [S9,S11], respectively. The past motion and future motions contain 25 frames (0.5 sec) and 100 frames (2.0 sec). AMASS unifies 24 different human motion datasets including HumanEva-I (sigal2010humaneva) with the SMPL (SMPL) pose representation. AMASS contains 9M frames at 60 Hz in total. As a multi-dataset collection of AMASS, one can perform a cross-dataset evaluation. We follow the evaluation protocol proposed by BeLFusion (barquero2023belfusion) for fair comparison, as predicting future 120 frames (2.0 sec) with 30 frames observation (0.5 sec) with downsampling to 60 Hz.
Metrics. We use the evaluation metrics to measure diversity and accuracy. 50 sampled predictions are evaluated with the following metrics: Average Pairwise Distance (APD) (aliakbarian2020stochastic) evaluates sample diversity. It calculates the mean $l_{2}$ distance between all predicted motions. Average and Final Displacement Error (ADE, FDE) (alahi2016social; lee2017desire; gupta2018social) evaluate accuracy. They calculate the average and final-frame $l_{2}$ distances between the ground truth motion and closest prediction in the 50 set. Multimodal ADE and FDE (MMADE, MMFDE) (yuan2019diverse) also evaluate accuracy in a similar way to ADE and FDE. However, they are calculated over multimodal ground truths selected by grouping similar motions.
@@
Implementation Details. Our method is based on a latent flow-based model as mentioned in Sec. 3.2. We utilize a Variational Autoencoder (VAE) for $\mathcal{E},\mathcal{D}$ to obtain a latent representation. Specifically, we employ the Behavioral Latent Space (BLS) (barquero2023belfusion) as a VAE to achieve a compact latent representation. BLS ensures smoothness of predicted motions and consistency between the end of the past motion and the start of the predicted motion. Additionally, we compress this representation using linear factorization (xu2024learning). The dimensionality of the VAE latent space is 128, which we further reduce to 8 dimensions through linear factorization. We trained the unconditional flow-based model on this 8-dimensional space. The unconditional flow-based model $f_{\theta}$ is a continuous normalizing flow (CNF) model, with its vector field regressed by a U-Net architecture. The conditional base density $q_{\phi}$, as well as the VAE encoder and decoder, are implemented as one-layer Recurrent Neural Networks (RNNs). We used a Gaussian mixture model with $M=50$ modes to model the conditional base density $q_{\phi}$. We precomputed and collected triplets $t_{k}=\{\bm{z}_{k},|{\rm det}\mathcal{J}_{f_{\theta}}(\bm{z}_{k})|^{-1},\bm{x}_{k}\}$ using all training samples of each dataset. All experiments, including inference time measuring, were carried out using a single NVIDIA A100 GPU. We used a batch size 64 and the Adam optimizer with a learning rate of $5\times 10^{-4}$.
@@ 5.1 Quantitative Evaluation
[…]
@@
The density estimation accuracy of each method is compared between CacheFlow and the state-of-the-art methods. The three state-of-the-art methods BeLFusion (barquero2023belfusion), CoMusion (sun2024comusion), and SLD (xu2024learning) are selected. CoMusion and SLD were selected since they outperform our method in benchmarks of stochastic human motion prediction. We also include BeLFusion to compare CacheFlow with the method with latent space. We applied KDE to these previous methods since they only sample a set of predictions and cannot estimate density. While we evaluated 50 and 1000 samples for KDE on BeLFusion and CoMusion, SLD only allows 50 samples due to the fixed number of anchors corresponding to predictions. We measured the inference time of each method to estimate the density of ten thousand future motions from the past motion input.
The quantitative comparisons over the multimodal ground truth log probability are shown in Tab.˜2. All previous methods suffer from slow inference of their own and KDE on high-dimensional motion data. Their inference time exceeded the prediction horizon of 2000ms in the future. Therefore, they cannot estimate density in real-time. In contrast, our method achieves better estimation accuracy in less than one millisecond. This indicates that CacheFlow has strong discriminative ability to list up possible future motions required for safety assurance. Our method is even faster only on the density estimation (0.5ms) than the inference time reported in Tab.˜1 (1.3ms). This is because we don’t need any extra sampling operation in the density estimation.
@@ 5.2 Qualitative Comparison of Predicted Motions
To visually evaluate CacheFlow, we conducted a qualitative comparison of methods on the AMASS dataset, as shown in Fig.˜3. We visualized the end poses of 10 samples from each method alongside the end poses of past motions and the ground truth future motions. The sitting or lying poses were translated to the ground plane, as the global translation is not modeled in human motion prediction. The 10 pose samples are arranged from the closest to the farthest from the ground truth pose based on joint rotations.
@@
Our observations indicate that CacheFlow predicts realistic poses. The closest poses to the ground truths also demonstrate that the accuracy of CacheFlow is comparable to CoMusion, as reflected in the ADE and FDE metrics listed in Tab.˜1. Notably, our method is computationally efficient, operating 100 times faster than the fastest CoMusion. In summary, CacheFlow effectively delivers realistic and accurate predictions.
5.3 Further Analysis
5.3.1 Generalization to Out-of-Distribution Data
Table 3: Zero-shot evaluation on 3DPW (3DPW). The models are trained on AMASS (mahmood2019amass). For CacheFlow, the cache are computed on AMASS training set also.
Zero-shot 3DPW (3DPW)
[…]
TPK (walker2017pose) 9.582 0.648 0.701 0.665 0.702
DLow (yuan2020dlow) 13.772 0.581 0.649 0.602 0.651
GSPS (mao2021generating) 11.809 0.552 0.650 0.578 0.653
DivSamp (dang2022diverse) 24.153 0.554 0.678 0.593 0.686
BeLFusion (barquero2023belfusion) 7.740 0.493 0.590 0.531 0.599
CoMusion (sun2024comusion) 11.404 0.477 0.570 0.540 0.587
SkeletonDiff.(curreli2025nonisotropic) 9.814 0.472 0.575 0.535 0.594
CacheFlow 7.963 0.499 0.591 0.539 0.601
To evaluate the generalization capability of CacheFlow on unseen domains, we conduct a zero-shot evaluation on the 3DPW dataset (3DPW). In this setting, all models are trained strictly on the AMASS dataset (mahmood2019amass) and tested directly on 3DPW without any fine-tuning. This poses a significant challenge due to the domain shift between the indoor optical motion capture data of AMASS and the in-the-wild outdoor sequences of 3DPW.
The results are summarized in Tab. 3. Despite the challenging zero-shot setting, CacheFlow demonstrates robust generalization performance. Specifically, our method achieves an ADE of 0.499 and an FDE of 0.591, surpassing earlier baselines such as TPK, DLow, and GSPS. Furthermore, our performance is highly competitive with recent diffusion-based approaches, such as BeLFusion (barquero2023belfusion) (0.493 ADE) and CoMusion (sun2024comusion) (0.477 ADE), showing only marginal differences in error metrics.
A crucial aspect of this experiment is the configuration of the cache. As noted in the table, the cache used by CacheFlow is constructed solely from the AMASS training set. The competitive performance on 3DPW indicates that CacheFlow does not simply memorize training samples. Instead, it successfully leverages the retrieved motion priors from the source domain (AMASS) to reconstruct valid motion trajectories in the target domain (3DPW). This suggests that the latent flow capabilities of CacheFlow effectively bridge the domain gap by generalizing the transitions found in the cached training data to unseen, in-the-wild scenarios.
CacheFlow generalizes beyond training data in two key ways. First, although cached latents $z_{k}$ are drawn from the training set, generated futures $\hat{X}$ differ significantly because they are conditioned on novel past contexts $c$. This allows the model to generate motions unseen during training. Second, the large scale of the Human3.6M (371,188 patterns) and AMASS (723,263 patterns) datasets already provides considerable diversity, and CacheFlow leverages this diversity to further expand to novel futures. Even with fixed cached codes, novel inputs yield generalized future motions that go beyond memorization of the training set.
5.3.2 Extensive Computational and Memory Cost Analysis
We conducted a detailed evaluation of both computational overhead and memory requirements for CacheFlow across motion prediction benchmarks, using an Intel(R) Xeon(R) Gold 6154 CPU and a single NVIDIA TITAN RTX GPU. The results demonstrate that our caching mechanism achieves efficient memory usage while maintaining practical inference times on both CPU and low-performance GPU platforms.
$\text{Memory Cost (MB)}=\text{Bytes per motion}\times\#\text{motion}\,/\,1024^{2}$
The measured costs are summarized in Table 4.
Table 4: Computational and memory cost analysis on Human3.6M and AMASS datasets.
Cost Memory Computation
Bytes per Mot. # Mot. Cache Size (MB) GPU mem. (MB) GPU CPU
Cache Inf. Cache Inf
Human3.6M 68 371188 24.07 811 26min 2.7ms 1h19min 3.6ms
AMASS 68 723263 46.90 875 1h10min 3.4ms 3h40min 6.9ms
This analysis highlights three important aspects. First, the cache footprint is very small, less than 50 MB for both datasets. Second, the preprocessing cost is modest, requiring only one to four hours even for AMASS on CPU. Third, the inference time is highly practical, amounting to only a few milliseconds per sample on both CPU and GPU, which makes CacheFlow suitable for real-time applications.
5.3.3 Ablation Study
[…]
We conducted an ablation study to investigate how each component affects the performance of our CacheFlow. We ablate five components: (1) dimensionality reduction via linear factorization on VAE, (2) the unconditional flow-based model $f_{\theta}$, (3) joint learning of the conditional base density $q_{\phi}$ and unconditional flow-based model $f_{\theta}$, (4) dataset for precomputation, (5) the sampling method for metrics over a fixed number of predictions. Ablation results on the Human3.6M dataset are summarized in Tab.˜5.
The Unconditional Flow-based Model. We ablate this flow-based model $f_{\theta}$ to confirm it improves the conditional base density $q_{\phi}$ by adding complexity. As shown in Tab.˜5, we observe a notable performance drop without the flow-based model. Therefore, our unconditional flow-based model $f_{\theta}$ complements conditional base density $q_{\phi}$ to estimate complex density distribution over human motions.
Cache Precomputation. We propose the precomputation over the training split. Specifically, we apply inverse transform $\bm{z}=f_{\theta}(\bm{x})$ to ground truth future motions in the training split. However, we may precompute infinite precomputation samples. For example, we can sample $\bm{z}\sim q_{\phi}(\bm{z}|\bm{c})$ and obtain $\bm{x}$ by forward transform $\bm{x}=f_{\theta}(\bm{z})$. As shown in the ablation, precomputation on the training split outperforms one on the base density since we can regularize the prediction to be legitimate human motions using the training split.
Sampling Method. We propose the nearest neighbor sampling from the precomputation set as described in Sec.˜4.3. Lastly, we ablate this sampling to evaluate its performance gain. We experimented with two sampling method alternatives: random sampling and most likely sampling. Precomputed motion features $\bm{x}_{k^{*}}$ are uniformly selected as predictions with random sampling. Most likely sampling selects motion features $\bm{x}_{k^{*}}$ with the highest probabilities $k^{*}={\rm argmax}_{k}p(\bm{x}_{k}|\bm{c})$. We found that the large and little performance drops with random and most likely sampling respectively. This random sampling is worse due to the independence from the past motions $\bm{c}$. The most likely method underperforms due to less diverse samples. It cannot select a motion feature set with diversity because all selected features are often located in one peak of the estimated density. Since ADE and FDE are best-of-many metrics, this less diversity leads to worse performance. In contrast, our sampling method is superior to others. Our sampling incorporates past motions and achieves good diversity by simulating sampling from the estimated density $p(\bm{x}|\bm{c})$.
@@ 6 Concluding Remarks
Appendix A Reproducibility Statement
We are committed to ensuring the reproducibility of our results. To this end, we will release the full source code, including training and evaluation scripts. All datasets used in this study (Human3.6M (ionescu2013h36m), AMASS (mahmood2019amass), and derived benchmarks) are publicly available. Our implementation is based on PyTorch, and all experiments except Sec.˜5.3.2 were conducted on a workstation with an Intel Xeon Gold 6154 CPU and an NVIDIA A100 GPU.
We report detailed computational and memory costs in Table 4, and we describe architectural and training hyperparameters in Appendix B. For reproducibility, we specify the number of training epochs, optimizer (Adam with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, learning rate $5\times 10^{-4}$), batch sizes (192 for Human3.6M, 64 for AMASS). Random seeds were fixed to 0 across all experiments.
Appendix B Implementation Details of Kernel Density Estimation